OpenStudy (anonymous):

5+(square root) 5 divided by 8- (square root) 5

OpenStudy (anonymous):

\[(5+\sqrt{5})/(8-\sqrt{5})\] Multiply both sides by (8+sqrt(5)) to rationalize the denominator\[(5+\sqrt{5})(8+\sqrt{5})/59\]\[= (35+13\sqrt{5})/59\]

OpenStudy (anonymous):

thank you